OpenStudy (anonymous):

The quantity demanded x for a product is inversely proportional to the cube of the price p for p > 1. When the price is $10 per unit, the quantity demanded is 125 units. The initial cost is $120 and the cost per unit is $4. What price will yield a maximum profit? (Round your answer to two decimal places.) I'm not sure how to do this, can somebody please help explain this to me?

OpenStudy (tkhunny):

"The quantity demanded x for a product is inversely proportional to the cube of the price " This should speak to you directly and say this: \(x = \dfrac{k}{p^{3}}\)

OpenStudy (anonymous):

x=k/(10^3)

OpenStudy (anonymous):

or would it be x=k/(120^3)

OpenStudy (tkhunny):

125 = k/(10^3) -- Solve for k

OpenStudy (anonymous):

soo would I multiply both sides by k? and then divide both by 125?

OpenStudy (tkhunny):

That makes no sense. Why not just multiply by 1000 and be done with it?

OpenStudy (anonymous):

ooh, yeah, that would be easier

OpenStudy (anonymous):

I have bad math dyslexia and mess up processes all the time, this is why i need help with simple algebra

OpenStudy (radar):

So what is the price that maximize profits. For amount shown 125*4 + 120 = $620 cost Gross received $1250.0 profit = $1250 - $620 = $630.00

OpenStudy (radar):

What is the problem asking of you??